### v1

@@ Appendix B Numerical Studies on GW, EW, and EWλ
Next, we compare randomly rotated gray-value images. We can describe such images as mm-spaces $\mathbb{X}_{i}=(X,d,\xi_{i})$ where $X\subset\mathbb{R}^{2}$ is a grid that describes the pixel positions equipped with the Euclidean metric and $\xi_{i}$ is the pixel intensity, see Beier et al. (2023). We use $(Z,d_{Z})=(X,d)$. We apply random affine transformations, i.e., translations and rotations, to the first 10 FashionMNIST Xiao et al. (2017) training images from the classes “Trouser”, “Pullover” and “Sneaker”, respectively. Then we compute $\operatorname{W}$, $\operatorname{GW}$ and $\operatorname{EW}_{\lambda}$ for $\lambda=20$. The results are displayed in Figure 21. As expected, the Wasserstein distance cannot recover the class structure due to the affine transformations, whereas $\operatorname{EW}_{\lambda}$ and GW capture the class structure. Moreover, we see that the distance matrices of $\operatorname{EW}_{\lambda}$ and GW display almost identical patterns, up to scaling.
@@
Appendix C Further Example for Comparing Feature Spaces
@@
As an addition to the genetyping experiment, we present another experiment with real-world data based on latent spaces. We consider the task of comparing the 4d latent spaces of an auto-encoder (AE) versus a variational auto-encoder (VAE), trained on FashionMNIST Xiao et al. (2017). For this purpose, we train a simple convolutional AE with two convolutional layers and one linear layer for the encoder and the decoder. We train the AE with the MSE loss and the VAE with the log-likelihood loss for 10 epochs on the canonical training data splits. Subsequently, we embed the test data using the resulting AE and VAE into the 4d latent spaces. Note that such distinct AE trained on the same dataset generally produce incomparable embeddings. For our experiment, we consider the first 100 data points of the test split. We choose $Z$ as an equispaced 20$\times$20 grid on $[0,1.3]\subset\mathbb{R}^{2}$ with the Euclidean distance. We follow the same evaluation routine as in Subsection 6.2 for parameter selection and visualization. The joint embeddings are visualized in Figure 24. Comparing FOSCTTM and KNN, we see again a good embedding quality for our approach, see the figure caption.

### v2

@@ Appendix B Numerical Studies on GW, EW, and EWλ
Next, we compare randomly rotated gray-value images. We can describe such images as mm-spaces $\mathbb{X}_{i}=(X,d,\xi_{i})$ where $X\subset\mathbb{R}^{2}$ is a grid that describes the pixel positions equipped with the Euclidean metric and $\xi_{i}$ is the pixel intensity, see (Beier et al., 2023). We use $(Z,d_{Z})=(X,d)$. We apply random affine transformations, i.e., translations and rotations, to the first 10 FashionMNIST (Xiao et al., 2017) training images from the classes “Trouser”, “Pullover” and “Sneaker”, respectively. Then we compute $\operatorname{W}$, $\operatorname{GW}$ and $\operatorname{EW}_{\lambda}$ for $\lambda=20$. The results are displayed in Figure 21. As expected, the Wasserstein distance cannot recover the class structure due to the affine transformations, whereas $\operatorname{EW}_{\lambda}$ and GW capture the class structure. Moreover, we see that the distance matrices of $\operatorname{EW}_{\lambda}$ and GW display almost identical patterns, up to scaling.
@@
Appendix C Parameter Sensitivity
Algorithm 1 depends on three input parameters, namely the GW regularization parameter $\lambda$, the entropic regularization parameter $\varepsilon$ and the chosen discrete reference space $Z$. Generally, we aim to set $\lambda$ as high as possible and $\varepsilon$ as low as possible while preserving numerical stability. Based on two sets of 3d shapes described in Section 6.1, namely a s-curve in combination with a Swiss roll with a hole and two human shapes, we present a parameter sensitivity study that illustrates the impact of $\lambda$, $\varepsilon$ and $Z$. Throughout this study note that our Euclidean joint embeddings are invariant to translations and rotations which results in arbitrarily oriented joint embeddings. In Figure 24, we investigate the influence of the GW parameter $\lambda$ that promotes near-isometric embeddings. We observe that the embeddings of both spaces nearly coincide for small $\lambda$. This is supported by Proposition 3.6 showing that the embeddings form a fixed-support barycenter in the limiting case. Moreover, we see a saturation effect where an increase of $\lambda$ from 1 to 100 has only a small impact on the joint embeddings. In Figure 27, we investigate the influence of the entropic parameter $\varepsilon$ that enables efficient computation. Here, we see that a large $\varepsilon$ leads to a highly blurred embedding. This shows that it is advisable to choose a low, but numerically stable value for $\varepsilon$. Lastly, we investigate the discretization of the reference space in Figure 30. Again, we aim for a fine discretization, but the runtime of our algorithm depends heavily on the size $|Z|$ of our reference space. We observe that the joint embeddings display the same shapes and alignments, only at higher resolution. All embeddings were computed with 40 iterations of Algorithm 1.
$\lambda=0.5$
$\lambda=1$
$\lambda=100$
(a) S-bended rectangle and Swiss roll with hole.
$\lambda=0.5$
$\lambda=1$
$\lambda=100$
(b) Same human shape in different poses.
Figure 24: Ablation study of $\lambda$ for joint embedding of shapes from Figure 8(b) and 11(a). Based on varying $\lambda$ values, we employ our method with $\varepsilon=0.001$ and $Z$ defined by a uniform $20\times 20$ grid in $[0,1.3]^{2}$.
$\varepsilon=0.001$
$\varepsilon=0.002$
$\varepsilon=0.01$
(a) S-bended rectangle and Swiss roll with hole.
$\varepsilon=0.001$
$\varepsilon=0.002$
$\varepsilon=0.01$
(b) Same human shape in different poses.
Figure 27: Ablation study of $\varepsilon$ for joint embedding of shapes from Figure 8(b) and 11(a). Based on varying $\varepsilon$ values, we employ our method with $\lambda=100$ and $Z$ defined by a uniform $20\times 20$ grid in $[0,1.3]^{2}$. Below the joint embeddings, we show convergence plots of the unregularized functional $F_{100}(\alpha)$ from (17) after 4, 8, 12, 16, 20, 24 and 28 iterations.
$|Z|=15^{2}$
$|Z|=20^{2}$
$|Z|=60^{2}$
(a) S-bended rectangle and Swiss roll with hole.
$|Z|=15^{2}$
$|Z|=20^{2}$
$|Z|=60^{2}$
(b) Same human shape in different poses.
Figure 30: Ablation study of the grid size for joint embedding of shapes from Figure 8(b) and 11(a). Based on $\varepsilon=0.001$ and $\lambda=100$, we employ our method with $Z\subset[0,1.3]^{2}$ defined by a uniform $15\times 15$, $20\times 20$ or $60\times 60$ grid. We include CPU runtimes.
Appendix D Hyperparameter Selection
JMDS, SCOT, and our method all rely on the Sinkhorn algorithm. Therefore, we use the same fixed entropic regularization $\varepsilon$ for all three methods for a given experiment. We choose it as small as possible while avoiding numerical overflow, i.e., $\varepsilon=0.001$ or $\varepsilon=0.002$. Next, there are multiple ways of defining the nearest neighbor graph described in Section 6.2. For our experiments, we start by running the SCOT algorithm which offers an unsupervised hyperparameter search to determine the shape of the nearest neighbor graph (Demetci et al., 2022). We use the unsupervised hyperparameter selection of the publicly available SCOT implementation [^3] https://github.com/rsinghlab/SCOT (March 2025) to estimate a suitable nearest-neighbor graph. Here, we use the same $\varepsilon$ as in the rest of the experiment and use the default grid over the neighborhood size. Note that the public implementation uses a connectivity graph based on pairwise correlation distances between points instead of Euclidean distances, see (Demetci et al., 2022). Then, we use SCOT and MDS to estimate the final embeddings based on the pairwise geodesic distance matrix estimated from this graph. We use the same distance graph for UnionCom, JMDS, and our algorithm. Here, we estimate suitable hyperparameters based on a random 10% validation split. We use four parameter configurations for each algorithm, i.e., we test $\lambda=0.01,0.1,10,100$ for JMDS and our method. For UnionCOM, we use combinations of $\beta=1,10$ and a perplexity of $30$ and $100$ for the t-SNE dimensionality reduction. We use no parameter annealing for any method.
Appendix E Further Example for Comparing Feature Spaces
@@
As an addition to the genetyping experiment, we present another experiment with real-world data based on latent spaces. We consider the task of comparing the 4d latent spaces of an auto-encoder (AE) versus a variational auto-encoder (VAE), trained on FashionMNIST (Xiao et al., 2017). For this purpose, we train a simple convolutional AE with two convolutional layers and one linear layer for the encoder and the decoder. We train the AE with the MSE loss and the VAE with the log-likelihood loss for 10 epochs on the canonical training data splits. Subsequently, we embed the test data using the resulting AE and VAE into the 4d latent spaces. Note that such distinct AE trained on the same dataset generally produce incomparable embeddings. For our experiment, we consider the first 100 data points of the test split. We choose $Z$ as an equispaced 20$\times$20 grid on $[0,1.3]\subset\mathbb{R}^{2}$ with the Euclidean distance. We follow the same evaluation routine as in Subsection 6.2 for parameter selection and visualization. The joint embeddings are visualized in Figure 33. Comparing FOSCTTM and KNN, we see again a good embedding quality for our approach, see the figure caption.